# ALERT Open Dataset and Input-Size-Agnostic Vision Transformer for Driver Activity Recognition using IR-UWB

### v1

@@ Abstract
Distracted driving contributes to fatal crashes worldwide. To address this, researchers are using driver activity recognition (DAR) with impulse radio ultra-wideband (IR-UWB) radar, which offers advantages such as interference resistance, low power consumption, and privacy preservation. However, two challenges limit its adoption: the lack of large-scale real-world UWB datasets covering diverse distracted driving behaviors, and the difficulty of adapting fixed-input Vision Transformers (ViTs) to UWB radar data with non-standard dimensions.
This work addresses both challenges. We present the ALERT dataset, which contains 10,220 radar samples of seven distracted driving activities collected in real driving conditions. We also propose the input-size-agnostic Vision Transformer (ISA-ViT), a framework designed for radar-based DAR. The proposed method resizes UWB data to meet ViT input requirements while preserving radar-specific information such as Doppler shifts and phase characteristics. By adjusting patch configurations and leveraging pre-trained positional embedding vectors (PEVs), ISA-ViT overcomes the limitations of naive resizing approaches. In addition, a domain fusion strategy combines range- and frequency-domain features to further improve classification performance.
Comprehensive experiments demonstrate that ISA-ViT achieves a 22.68% accuracy improvement over an existing ViT-based approach for UWB-based DAR. By publicly releasing the ALERT dataset and detailing our input-size-agnostic strategy, this work facilitates the development of more robust and scalable distracted driving detection systems for real-world deployment.
@@ I Introduction
According to the World Health Organization (WHO), 1.19 million people die every year due to vehicle accidents [who]. Additionally, the National Highway Traffic Safety Administration (NHTSA) [NHTSA] in the United States estimates that approximately 6 million traffic accidents are reported to police each year, resulting in 38,000 to 40,000 fatalities annually. Among these vehicle accidents, distracted driving accounts for about 10% of fatal incidents, often involving some form of driver distraction. Given its unavoidable nature, detecting and mitigating distracted driving is a crucial task.
@@
To prevent driver distraction, many researchers have studied Driver Activity Recognition (DAR) using various approaches: camera vision [ohn2014head, kavi2016multiview], acoustic signals [xie2021hearsmoking, xie2019d], and Radio Frequency (RF) signals [wifiheadtracking, jia2018wifind, bai2019widrive, bai2020carin, RaDA, blinkradar, wificamera]. However, these approaches have several limitations. To prevent driver distraction, many researchers have studied Driver Activity Recognition (DAR) using various approaches: camera vision [ohn2014head, kavi2016multiview], acoustic signals [xie2021hearsmoking, xie2019d], and Radio Frequency (RF) signals [wifiheadtracking, jia2018wifind, bai2019widrive, bai2020carin, RaDA, blinkradar, wificamera]. However, these approaches have several limitations. Camera vision struggles with poor lighting and raises privacy concerns regarding the potential restoration in intelligent vehicle systems [Hwang_2023_ICCV, concretizer]. Acoustic signal-based methods can be disrupted by ambient noise and also pose privacy risks due to continuous microphone activation. RF signal-based approaches have predominantly been studied using WiFi. However, utilizing WiFi, primarily operating within the busy 2.4 GHz industrial, scientific, and medical (ISM) band, within a vehicle can potentially cause interference with other vehicle devices using Bluetooth or similar busy ISM band communications.
@@
In this paper, we exploit the Impulse Radio Ultra-Wideband (IR-UWB) radar for DAR while driving to tackle these problems. UWB offers several advantages: i) robustness against interference and multipath fading due to its wide bandwidth ($\geq$ 500 MHz), ii) operation within the 6.5–9 GHz range with low transmit power (-41.3 dBm/MHz), ensuring coexistence with popular wireless formats such as satellite navigation, WiFi, Bluetooth, and ZigBee [fira]. iii) inherent privacy protection, as it avoids capturing visual or audio data. Consequently, UWB-based Human Activity Recognition (HAR) has garnered significant attention [HARSANET, HAR_2017, SleepPoseNet, RaDA, blinkradar, Uhead, han2020ir, maitre2021recognizing, noori2021ultra, maitre2020fall, vit1, vit2].
@@
Challenges. Despite the potential of UWB, applying UWB to DAR presents two key challenges. The first is the lack of comprehensive research addressing diverse and complex distracted driving activities using UWB sensing data in real-world driving environments. Collecting such datasets demands substantial time and effort, with the workload increasing proportionally to the number of labels and environmental complexity. This difficulty may partly explain why many studies [xie2021hearsmoking, xie2019d, Uhead, blinkradar] have focused solely on single distracted driving activities. However, given the varied nature of distracted driving behaviors, monitoring only one type of behavior is insufficient to prevent traffic accidents.
@@
Although existing DAR dataset [RaDA] have collected dataset on various distracted driving activities, the dataset is gathered in simulated environments. Such an approach fails to account for critical real-world factors, such as road conditions, vehicle vibrations, and other environmental influences. This limitation raises concerns about the generalizability and reliability of these datasets for real driving scenarios.
The second challenge lies in the incompatibility of UWB data with state-of-the-art models for DAR. Recently, Vision Transformers (ViTs) [vit, han2022survey, khan2022transformers, ruan2022vision] have emerged as state-of-the-art models across multiple vision task benchmarks, gaining significant attention for their performance in classification tasks. Consequently, many researchers have attempted to transfer the powerful capabilities of ViTs to other domains, including UWB domain. However, resizing UWB radar images to align with standard ViT input dimensions is non-trivial, as naive interpolation can distort essential radar-specific information such as Doppler shifts, phase, and signal attenuation. Pre-trained ViTs rely on positional embedding vectors (PEVs) optimized for fixed-size natural images, and mismatches in spatial relationships degrade performance.
@@
Nevertheless, many studies [li2023human, zhao2022distributed, lai2023vision, li2023advancing, kim2023study] using ViTs for UWB sensing have treated this resizing step as a minor implementation detail, not detailing any domain-specific approach to handle the unique geometry of UWB data. This oversight can result in significant performance loss, highlighting the need for an input-size-agnostic approach that preserves crucial radar-domain features while seamlessly integrating with pre-trained ViTs.
@@
Approaches. To address the first challenge, the lack of real-driving datasets, we collect the ALERT dataset, comprising 10,220 samples of 7 distracted driving activities captured in real-driving environments. We present benchmarking performance of widely used 8 learning algorithms—CNN-based [googlenet, resnet, densenet, mobilenet], RNN-based [maitre2020fall, maitre2021recognizing, noori2021ultra], and transformer-based approaches [deit, vit]—that are applicable in UWB DAR. Additionally, we design experiments to provide the effective utilization of the ALERT dataset across these 8 algorithms such as analyzing observation time, multipath effects, and frequency area. Consequently, we make the ALERT dataset publicly available to promote its adoption and foster further advancements in UWB DAR research.
@@ II Related Works
DAR using UWB radar. Recently, several DAR systems utilizing UWB technology have emerged [blinkradar, RaDA, Uhead, khan2022ir]. Xu et al[Uhead] present UHead, a driver attention monitoring system based on UWB radar. Unlike traditional methods that rely on specialized sensors or cameras, UHead tracks the driver’s head motion to infer attention. It extracts time-frequency data from UWB radar signals to estimate the direction and angle of head rotations. However, detecting a driver’s distracted behavior solely based on head rotation has its limitations. Drivers can exhibit distracted behaviors without rotating their heads significantly such as using a mobile phone or smoking while driving.
Similarly, Hu et al. [blinkradar] design a system for fine-grained eye-blink monitoring in driving scenarios using customized IR-UWB radar. The system extracts eye-blink signals while filtering out interference from body movements and environmental factors. However, this approach is limited as it only targets drowsiness, overlooking other potential distractions that may occur during driving.
Khan et al. [khan2022ir] present a radar-based system for robust heart rate detection intended for in-car monitoring. They propose a novel time-series-based algorithm instead of traditional frequency-domain methods, using a deep learning classifier to detect heart rate patterns. Additionally, they optimize the radar sensor’s location inside the car to reduce interference from random body motions.
@@
The study in [RaDA] introduces an IR-UWB radar-based monitoring system capable of detecting a wide range of in-car activities. It also provides an open dataset for DAR in a simulated driving environment, covering 6 distracted driving activities: driving, autopilot, sleeping, driving while using a smartphone, smartphone use, and talking to a passenger. To the best of our knowledge, this is the first open dataset for UWB DAR in a simulated driving environment addressing six distracted driving activities.
However, simulated-driving environments fail to accurately represent real-world driving conditions, which involve factors such as ground vibrations, vehicle vibrations, and driving realism. For example, Morales-Alvarez et al[morales] report that the recognition accuracy of driver observation models decreased from 85.7% to 46.6% when transferred from simulation to real-world scenarios. Similarly, Stocco et al. [stocco] emphasize the “reality gap” in simulation-based testing for autonomous systems, noting that critical cues like vibration, cabin constraints, and natural driver reactions are often missing in simulated settings.
@@
Compared to the dataset from [RaDA], the ALERT dataset focuses on more detailed distracted driving activities in real-world driving environments. It includes data for 7 activities: relaxation (autopilot), driving, nodding, smoking, drinking, panel control, and smartphone use. Additionally, the ALERT dataset provides both range and frequency domain data, enabling a variety of experimental approaches to be explored using the data.
The use of ViT with UWB. Various works [li2023human, zhao2022distributed, lai2023vision, li2023advancing, kim2023study] introduce the use of ViT with UWB. Li et al[li2023human] introduce MobileViTX, an improved version of the MobileViT architecture, specifically designed for human activity recognition (HAR) using IR-UWB radar. The Zhao et al[zhao2022distributed] explore the feasibility of classifying human activities using data captured by a distributed UWB radar system. They compare ViT with CNN-based architectures and validate the robustness of these models, demonstrating ViT’s ability to generalize to unseen participants.
Moreover, Lai et al. [lai2023vision] focus on developing a non-invasive sleep posture recognition system using multiple UWB radars. They employ machine learning models for classification, including CNN-based networks and ViT. Kim et al[kim2023study] introduce a new approach for 3D human pose estimation using IR-UWB radar, leveraging a transformer-based deep learning model.
The above studies have attempted to apply UWB data to ViT. However, none of them provide sufficient discussion on how to adapt UWB data, which has a different input size from images, for use in ViT. The unique study [ast] addresses how to apply audio data to ViT when the input shapes differ. Gong et al[ast] introduce the audio spectrogram transformer (AST) model, the first purely attention-based model for audio classification. They cut and interpolate the PEVs for height or width in the AST model. By doing so, they transfer 2D spatial knowledge from the pre-trained ViT to the AST model, even when the input shapes are different. However, excessive cutting or interpolation of PEVs may incur the ruin of spatial information in pre-trained ViT [ruan2022vision]. Therefore, we need to explore a method that is not affected by input size while still utilizing the pre-trained parameters of the ViT.
@@ III-A Sensor Specification
We capture the ALERT dataset using the UWB radar sensor Xethru X4M06, developed by Novelda [x4m06], as depicted in Fig. II. The X4 series are the off-the-shelf devices in UWB radar technology and are widely utilized in several UWB HAR studies [RaDA, HARSANET, SleepPoseNet]. The radar features two directional patch antennas (Tx and Rx), optimized for frequencies between 7.25–10.2 GHz with an opening angle of 65 in both azimuth and elevation. The detailed parameters of the radar are presented in Table II.
@@ III-C Data Collection
Our dataset exhibits a significant strength: We collect the dataset in a real-driving environment. Accordingly, we mount the radar within an actual vehicle, as depicted in Fig. 3(a). We strategically attach the radar to the vehicle’s air vent. This position does not obstruct the driver’s view while driving and is well-situated to capture the driver’s upper body and movements. Moreover, considering that modern smartphones are already equipped with built-in UWB chipsets, we have explored the potential integration of UWB radar into mobile phones. Since most drivers already use smartphone holders clipped to the air vent, mounting the UWB radar in the same location is not only practical but also aligns well with real-world usage patterns, supporting easy installation and potential commercialization.
@@
We have driven along the designated red line and collected the dataset, as shown in Figs. 3(b) and 3(c). The urban route spans 12 km and provides relatively smooth driving conditions, similar to highway driving, due to light traffic and few traffic signals. The route is predominantly flat, with a single uphill and downhill section, both having an approximate gradient of 10 degrees. Most of the road surface is paved with asphalt, although some areas are made of tiled-concrete as shown in Fig. 4(a) and 4(b). As a result, vehicle vibrations remain generally stable and are primarily influenced by the irregularities of the road surface.
The campus route forms a 6 km loop with a speed limit of 30 km/h, enabling low-speed driving. Due to the presence of frequent pedestrian crossings, vehicles often experience stop-and-go movement, creating conditions similar to traffic congestion. The route is composed of approximately 50% sloped sections—both inclines and declines—and 50% flat terrain. The average gradient of the uphill segments is around 10.2 degrees, while the downhill sections average about 9.2 degrees. Although the majority of the road surface is asphalt, the route also includes cobblestone areas, dirt roads, and numerous speed bumps as shown in Fig. 4(c)4(f), leading to diverse and irregular vibrations originating from the road surface.
@@
To detect the driver’s activity while driving, we select seven labels: Relaxation (Relax), steering wheel control (Drive), nodding (Nod), smoking (Smoke), drinking (Drink), panel control (Panel), and using smartphone (Phone). Except for Drive, the drivers must not perform all other activities while driving. According to the National Highway Traffic Safety Administration (NHTSA) [NHTSA], the Governors Highway Safety Association (GHSA) [GHSA], the Centers for Disease Control and Prevention (CDC) [CDC], and the Insurance Institute for Highway Safety (IIHS) [IIHS], National Institutes of Health (NIH) [NIH] in the United States and World Health Organization (WHO) [WHO_phone], five of these activities—Nod, Smoke, Drink, Panel, and Phone—are considered distracted driving and cause severe traffic accidents commonly observed in daily life.
Additionally, we include the Relax label to address the current issue of blind trust in autonomous driving technology. Many drivers using autonomous driving technology tend to disengage from driving and take their hands off the steering wheel. This behavior poses significant risks, highlighting the need for monitoring even seemingly benign activities like relaxation while driving. We explain seven activities in detail as follows.
  • Relax refers to a posture where drivers remove their hands from the steering wheel to rest, relying on autonomous driving technology. The drivers take breaks in various ways, such as freely taking their hands off the steering wheel and folding their arms, humming along to music, or talking on the phone using hands-free devices.
  • Drive involves drivers placing both hands on the steering wheel and driving through the courses shown in Fig. 3. The drivers are free to make left turns, right turns, or U-turns along the course according to their driving habits.
  • Nod is a key indicator of drowsy driving. The drivers nod their heads while keeping their hands on the steering wheel.
  • Smoke denotes smoking activities. The drivers freely smoke regular cigarettes and electronic cigarettes according to individual smoking habits. This activity involves inhaling and exhaling smoke, as well as putting ashes into the ashtray located in front of the armrest.
  • Drink involves drinking activities, either directly from a bottle or using a straw. The drivers place the bottle or cup in the cup holder, drink water or other beverages, and then return it to the cup holder located in front of the armrest.
  • Panel encompasses various tasks such as adjusting the stereo volume, air conditioning, controlling the navigation, and pressing panel buttons.
  • Phone involves various smartphone activities. The drivers exchange messages, swipe display to search for music playlist, surfs web, etc.
@@
We have obtained the dataset from nine volunteers, yielding a total of 10,220 samples (per 5 seconds) as shown in Table III. The participants consist of individuals with heights ranging from 165 to 188 cm and weights between 53 and 100 kg. Their ages range from 26 to 35 years old, with an average age of 32. Their driving experience spans from 1 to 10 years, with an average of 6 years. All participants are familiar with typical passenger vehicles, primarily sedans and SUVs. In addition, all participants reported prior experience with the types of distracted driving activities included in the ALERT dataset.

### v2

@@ Abstract
Distracted driving contributes to fatal crashes worldwide. To address this, researchers are using Driver Activity Recognition (DAR) with Impulse Radio Ultra-Wideband (IR-UWB) radar, which offers advantages like interference resistance, low-power use, and privacy. However, two challenges limit its adoption: the lack of large-scale, real-world UWB datasets on diverse distracted driving behaviors, and the difficulty in adapting fixed-input Vision Transformers (ViTs) to UWB radar data with non-standard dimensions.
This work tackles both challenges. We present the ALERT dataset, containing 10,220 radar samples of seven distracted driving activities in real driving conditions. We also propose the Input-Size-Agnostic Vision Transformer (ISA-ViT), a framework designed for radar-based DAR. ISA-ViT resizes UWB data to fit ViT input requirements while preserving radar-specific information like Doppler shifts and phase data. By adjusting patches and using pre-trained positional embedding vectors (PEVs), ISA-ViT avoids the limitations of simple resizing. Additionally, a domain fusion strategy combines range and frequency domain features, enhancing classification accuracy.
Comprehensive experiments demonstrate that ISA-ViT achieves a 22.68% higher accuracy compared to the existing ViT method in UWB-based DAR. By publicly releasing the ALERT dataset and detailing our input-size-agnostic strategy, this work paves the way for more robust and scalable distracted driving detection systems in real-world scenarios.
@@ I Introduction
According to the World Health Organization (WHO), 1.19 million people die every year due to vehicle accidents [41]. Additionally, the National Highway Traffic Safety Administration (NHTSA) [36] in the United States estimates that approximately 6 million traffic accidents are reported to police each year, resulting in 38,000 to 40,000 fatalities annually. Among these vehicle accidents, distracted driving accounts for about 10% of fatal incidents, often involving some form of driver distraction. Given its unavoidable nature, detecting and mitigating distracted driving is a crucial task.
@@
To prevent driver distraction, many researchers have studied Driver Activity Recognition (DAR) using various approaches: camera vision [40, 24], acoustic signals [51, 53], and Radio Frequency (RF) signals [52, 23, 2, 4, 6, 19, 3]. However, these approaches have several limitations. Camera vision struggles with poor lighting and raises privacy concerns regarding the potential restoration in intelligent vehicle systems [21, 28]. Acoustic signal-based methods can be disrupted by ambient noise and also pose privacy risks due to continuous microphone activation. RF signal-based approaches have predominantly been studied using WiFi. However, utilizing WiFi, primarily operating within the busy 2.4 GHz industrial, scientific, and medical (ISM) band, within a vehicle can potentially cause interference with other vehicle devices using Bluetooth or similar busy ISM band communications.
@@
In this paper, we exploit the Impulse Radio Ultra-Wideband (IR-UWB) radar for DAR while driving to tackle these problems. UWB offers several advantages: i) robustness against interference and multipath fading due to its wide bandwidth ($\geq$ 500 MHz), ii) operation within the 6.5–9 GHz range with low transmit power (-41.3 dBm/MHz), ensuring coexistence with popular wireless formats such as satellite navigation, WiFi, Bluetooth, and ZigBee [12]. iii) inherent privacy protection, as it avoids capturing visual or audio data. Owing to these characteristics, UWB sensing has been widely adopted for human activity recognition in indoor environments, where stable and robust sensing performance has been consistently reported [1]. Consequently, UWB-based Human Activity Recognition (HAR) has garnered significant attention [9, 44, 42, 6, 19, 54, 15, 34, 38, 33, 32, 55].
@@
Challenges. Despite the potential of UWB, applying it to DAR involves two key challenges. The first is the lack of comprehensive studies that address diverse and complex distracted driving activities using UWB sensing data collected in real-driving environments. Building such datasets is time-consuming and labor-intensive, and the effort grows with the number of labels and the complexity of driving conditions. This burden may partly explain why many studies [51, 53, 54, 19] focus on a single distracted activity. However, given the wide variety of distracted behaviors, monitoring only one activity is insufficient for preventing traffic accidents.
@@
Although an existing DAR dataset [6] includes multiple distracted driving activities, it is collected in simulated environments. As a result, it does not fully reflect real-world factors such as road conditions, vehicle-induced vibrations, and other environmental influences, which can limit generalizability to real driving scenarios.
The second challenge is the mismatch between UWB data and state-of-the-art DAR models. Vision Transformers (ViTs) [11, 16, 26, 43] have achieved strong performance across vision benchmarks and have therefore been adopted in UWB sensing. However, aligning UWB representations with fixed ViT input sizes is non-trivial: naive interpolation can distort radar-specific cues (e.g., Doppler shifts, phase, and attenuation), and pre-trained positional embedding vectors (PEVs) optimized for natural images may no longer match the UWB geometry, degrading performance.
@@
Nevertheless, many ViT-based UWB studies [31, 56, 29, 30, 27] treat resizing as a minor implementation detail and do not provide domain-specific handling of UWB input geometry. This omission can cause substantial performance loss, motivating an input-size-agnostic strategy that preserves radar-domain features while remaining compatible with pre-trained ViTs.
Research questions. To address these challenges, this study investigates the following research questions:
  • Can IR-UWB radar sensing reliably recognize diverse distracted driving activities in real-world driving environments?
  • How can Vision Transformers be effectively adapted to UWB radar data with varying input sizes without distorting domain-specific information?
  • Does fusing range and frequency domain representations improve the performance of UWB-based driver activity recognition?
@@
Approaches. To address the first challenge, the lack of real-driving datasets, we collect the ALERT dataset, comprising 10,220 samples of 7 distracted driving activities captured in real-driving environments. We present benchmarking performance of widely used 8 learning algorithms—CNN-based [46, 17, 20, 18], RNN-based [33, 34, 38], and transformer-based approaches [47, 11]—that are applicable in UWB DAR. Additionally, we design experiments to provide the effective utilization of the ALERT dataset across these 8 algorithms such as analyzing observation time, multipath effects, and frequency area. Consequently, we make the ALERT dataset publicly available to promote its adoption and foster further advancements in UWB DAR research.
@@
The rest of this paper is organized as follows. Sec. II reviews related works on IR-UWB-based activity recognition and vision-transformer-based modeling. Sec. III introduces the ALERT dataset and the preprocessing to construct range–time and frequency–time representations. Sec. IV describes the experimental protocol, baseline configurations, and evaluation metrics for fair benchmarking. Sec. V presents the proposed input-size-agnostic ViT with domain fusion. Sec. VI reports the experimental results and analyses. Sec. VII discusses key observations and limitations, and Sec. VIII concludes the paper.
@@ II Related Works
DAR using UWB radar. Recently, several DAR systems utilizing UWB technology have emerged [19, 6, 54, 25]. Xu et al[54] present UHead, a driver attention monitoring system based on UWB radar. Unlike traditional methods that rely on specialized sensors or cameras, UHead tracks the driver’s head motion to infer attention. It extracts time-frequency data from UWB radar signals to estimate the direction and angle of head rotations. However, detecting a driver’s distracted behavior solely based on head rotation has its limitations. Drivers can exhibit distracted behaviors without rotating their heads significantly such as using a mobile phone or smoking while driving.
Similarly, Hu et al. [19] design a system for fine-grained eye-blink monitoring in driving scenarios using customized IR-UWB radar. The system extracts eye-blink signals while filtering out interference from body movements and environmental factors. However, this approach is limited as it only targets drowsiness, overlooking other potential distractions that may occur during driving.
Khan et al. [25] present a radar-based system for robust heart rate detection intended for in-car monitoring. They propose a novel time-series-based algorithm instead of traditional frequency-domain methods, using a deep learning classifier to detect heart rate patterns. Additionally, they optimize the radar sensor’s location inside the car to reduce interference from random body motions.
@@
The study in [6] introduces an IR-UWB radar-based monitoring system capable of detecting a wide range of in-car activities. It also provides an open dataset for DAR in a simulated driving environment, covering 6 distracted driving activities: driving, autopilot, sleeping, driving while using a smartphone, smartphone use, and talking to a passenger. To the best of our knowledge, this is the first open dataset for UWB DAR in a simulated driving environment addressing six distracted driving activities.
However, simulated-driving environments fail to accurately represent real-world driving conditions, which involve factors such as ground vibrations, vehicle vibrations, and driving realism. For example, Morales-Alvarez et al[35] report that the recognition accuracy of driver observation models decreased from 85.7% to 46.6% when transferred from simulation to real-world scenarios. Similarly, Stocco et al. [45] emphasize the “reality gap” in simulation-based testing for autonomous systems, noting that critical cues like vibration, cabin constraints, and natural driver reactions are often missing in simulated settings.
@@
Compared to the dataset from [6], the ALERT dataset focuses on more detailed distracted driving activities in real-world driving environments. It includes data for 7 activities: relaxation (autopilot), driving, nodding, smoking, drinking, panel control, and smartphone use. Additionally, the ALERT dataset provides both range and frequency domain data, enabling a variety of experimental approaches to be explored using the data.
The use of ViT with UWB. Various works [31, 56, 29, 30, 27] introduce the use of ViT with UWB. Li et al[31] introduce MobileViTX, an improved version of the MobileViT architecture, specifically designed for human activity recognition (HAR) using IR-UWB radar. The Zhao et al[56] explore the feasibility of classifying human activities using data captured by a distributed UWB radar system. They compare ViT with CNN-based architectures and validate the robustness of these models, demonstrating ViT’s ability to generalize to unseen participants.
Moreover, Lai et al. [29] focus on developing a non-invasive sleep posture recognition system using multiple UWB radars. They employ machine learning models for classification, including CNN-based networks and ViT. Kim et al[27] introduce a new approach for 3D human pose estimation using IR-UWB radar, leveraging a transformer-based deep learning model.
The above studies have attempted to apply UWB data to ViT. However, none of them provide sufficient discussion on how to adapt UWB data, which has a different input size from images, for use in ViT. The unique study [14] addresses how to apply audio data to ViT when the input shapes differ. Gong et al[14] introduce the audio spectrogram transformer (AST) model, the first purely attention-based model for audio classification. They cut and interpolate the PEVs for height or width in the AST model. By doing so, they transfer 2D spatial knowledge from the pre-trained ViT to the AST model, even when the input shapes are different. Beyond vision and audio, transformer-based models have also been explored for non-visual spatio-temporal and time-series datasets, including 2D transformer formulations for financial time-series prediction and temporal graph-based modeling that captures structural and temporal dependencies [8, 48, 50]. However, excessive cutting or interpolation of PEVs may incur the ruin of spatial information in pre-trained ViT [43]. Therefore, we need to explore a method that is not affected by input size while still utilizing the pre-trained parameters of the ViT.
@@ III-A Sensor Specification
We capture the ALERT dataset using the UWB radar sensor Xethru X4M06, developed by Novelda [39], as depicted in Fig. II. The X4 series are the off-the-shelf devices in UWB radar technology and are widely utilized in several UWB HAR studies [6, 9, 42]. The radar features two directional patch antennas (Tx and Rx), optimized for frequencies between 7.25–10.2 GHz with an opening angle of 65 in both azimuth and elevation. The detailed parameters of the radar are presented in Table II.
@@ III-C Data Collection
Our dataset exhibits a significant strength: We collect the dataset in a real-driving environment. Accordingly, we mount the radar within an actual vehicle, as depicted in Fig. 3(a). We strategically attach the radar to the vehicle’s air vent. This position does not obstruct the driver’s field of view and is located at a height similar to the driver’s upper body (eye and chest level), enabling effective capture of driver movements with minimal occlusion from passengers or interior structures. By fixing the radar at this location, a stable and repeatable sensing geometry is maintained during driving, which is essential for consistent signal acquisition.
Moreover, considering that modern smartphones are already equipped with built-in UWB chipsets, we have explored the potential integration of UWB radar into mobile phones. Since most drivers already use smartphone holders clipped to the air vent, mounting the UWB radar in the same location is not only practical but also aligns well with real-world usage patterns, supporting easy installation and potential commercialization.
@@
[…]
Together, the two routes cover contrasting real-driving conditions—smooth and steady motion versus low-speed stop-and-go with frequent slope changes and road-surface transitions—thus providing diverse acquisition scenarios while maintaining safe and repeatable data collection.
@@
To detect the driver’s activity while driving, we select seven labels: Relaxation (Relax), steering wheel control (Drive), nodding (Nod), smoking (Smoke), drinking (Drink), panel control (Panel), and using smartphone (Phone). Except for Drive, the drivers must not perform all other activities while driving. According to the National Highway Traffic Safety Administration (NHTSA) [36], the Governors Highway Safety Association (GHSA) [13], the Centers for Disease Control and Prevention (CDC) [7], and the Insurance Institute for Highway Safety (IIHS) [22], National Institutes of Health (NIH) [37] in the United States and World Health Organization (WHO) [49], five of these activities—Nod, Smoke, Drink, Panel, and Phone—are considered distracted driving and cause severe traffic accidents commonly observed in daily life.
Additionally, we include the Relax label to address the current issue of blind trust in autonomous driving technology. Many drivers using autonomous driving technology tend to disengage from driving and take their hands off the steering wheel. This behavior poses significant risks, highlighting the need for monitoring even seemingly benign activities like relaxation while driving. We explain seven activities as follows: Relax: the driver keeps a relaxed posture with hands off the wheel; Drive: normal driving with both hands on the wheel; Nod: drowsiness-like head dropping and recovery; Smoke: bringing a cigarette to the mouth and lowering the hand; Drink: lifting a drink to the mouth and returning it; Panel: reaching toward and interacting with the center console (e.g., buttons/controls); Phone: raising a phone to the ear (or near the head) and lowering it.
@@
We have obtained the dataset from nine volunteers  [^1] All data were collected with informed consent and in accordance with applicable ethical standards. The UWB radar signals were anonymized and do not contain visual, audio, or personally identifiable information, thereby preserving participants’ privacy even for potentially sensitive behaviors., yielding a total of 10,220 samples (per 5 seconds) as shown in Table III. The participants consist of individuals with heights ranging from 165 to 188 cm and weights between 53 and 100 kg. Their ages range from 26 to 35 years old, with an average age of 32. Their driving experience spans from 1 to 10 years, with an average of 6 years. All participants are familiar with typical passenger vehicles, primarily sedans and SUVs. In addition, all participants reported prior experience with the types of distracted driving activities included in the ALERT dataset.